### v2

# Evolving Symbolic 3D Visual Grounder with Weakly Supervised Reflection

Abstract
3D visual grounding (3DVG) is challenging because of the requirement of understanding on visual information, language and spatial relationships. While supervised approaches have achieved superior performance, they are constrained by the scarcity and high cost of 3D vision-language datasets. On the other hand, LLM/VLM based agents are proposed for 3DVG, eliminating the need for training data. However, these methods incur prohibitive time and token costs during inference. To address the challenges, we introduce a novel training-free symbolic framework for 3D visual grounding, namely Evolvable Symbolic Visual Grounder (EaSe ), that offers significantly reduced inference costs compared to previous agent-based methods while maintaining comparable performance. EaSe uses LLM generated codes to compute on spatial relationships. EaSe also implements an automatic pipeline to evaluate and optimize the quality of these codes and integrate VLMs to assist in the grounding process. Experimental results demonstrate that EaSe achieves 52.9% accuracy on Nr3D dataset and 49.2% Acc@0.25 on ScanRefer, which is top-tier among training-free methods. Moreover, it substantially reduces the inference time and cost, offering a balanced trade-off between performance and efficiency. Codes are available at https://github.com/OpenRobotLab/EaSe.
keywords:
3D Visual Grounding, Large Language Model, Code Generation
[…]
figures/teaser.pdf Comparision between two previous training-free 3DVG methods and our method(EaSe ). For a query, agent based methods employ multimodal LLM to process scene information. They are more accurate, but their online LLM generation increases time consumption. Visual programming (Visprog.) method uses offline annoated relation functions, thus reduces grounding time, but it doesn’t perform well. In contrast, EaSe utilizes offline LLM generation and optimization before grounding and improves relation functions to relation encoders. As a result, EaSe ’s accuracy is close to agents but it’s consumption is much lower.
1 1. Introduction
The 3D visual grounding (3DVG) task focuses on locating an objects in a 3D scene based on a referring utterance. Numerous supervised methods have been proposed for 3DVG (Hsu et al., 2023; Jain et al., 2022; Huang et al., 2022; Chen et al., 2022; Huang et al., 2024; Ziyu et al., 2023; Bakr et al., 2023; Wu et al., 2023). These methods learn representations of referring utterance, object attributes and spatial relations from large-scale 3D vision language training data with high-quality annotations and achieve SOTA performances on 3DVG. However, 3D vision language datasets (Chen et al., 2020; Achlioptas et al., 2020) are scarce and their annotations are expensive. Additionally, some supervised methods are trained on these close-vocabulary datasets, which restricts their applicability in open-vocabulary scenarios. In recent years, large language models (LLMs) and vision-language models (VLMs) have shown remarkable capabilities in reasoning, code generation and visual perception. Building on these advancements, open-vocabulary and zero shot 3DVG agents (Yang et al., 2023; Xu et al., ; Fang et al., 2024; Li et al., 2024a) are proposed. These methods let LLM perform numerical computing and reasoning on object locations in the text modality (Yang et al., 2023; Fang et al., 2024), or let a VLM locate targets from scene scan images in the visual modality (Xu et al., ). Their responses which contain planning, reasoning or self-debugging processes are usually long. However, the generation is online, which means for every grounding query, the LLM generates long responses. As a result, they have significant costs in terms of grounding time and token usage (see Agents block in LABEL:fig:banner).
In contrast, visual programming method (Yuan et al., 2024b) utilizes LLM to generate a program which uses annotated relation functions and outputs the target object by executing the program. The generated program is short so its time and token consumption is much lower. However, it has trouble considering many spatial relations in the referring utterance simultaneously (Yuan et al., 2024a). This causes its relatively low accuracy (see Visprog. in LABEL:fig:banner).
To the end, we introduce EaSe, a novel training-free symbolic framework integrates LLM and VLM for 3D visual grounding which balances accuracy and inference cost. EaSe derives from previous neuro-symbolic frameworks (Hsu et al., 2023; Feng et al., 2024) and use LLM generated and optimized Python codes as spatial relation encoders (see  Figure 1, block (b)). EaSe also employs a VLM to process visual information of objects, which is also important but can’t be handled by code.
Specifically, EaSe parses the referring utterance into a symbolic expression which encapsulates all mentioned object categories and their spatial relations. Given positions of all objects in the scene, the spatial relation encoders generate relation features which can represent spatial relations between them. Then a special executor aggregates the symbolic expression, relation features and object categories to exclude most of objects that don’t match the referring utterance. Finally, the VLM selects the target from images of remaining candidate objects (see Figure 1, block (c)).
For more accurate relation encoders, we generate them through iterative optimization processes instead of straightly prompting the LLM. We introduce test suites to evaluate relation encoders. Not only the test suites enable us to choose better relation encoders from LLM responses, but also the LLM can leverage failed test cases to optimize its responses. Compared to the online generation of agent based methods, our generation and optimization is offline. Our relation encoders are reused in all grounding processes, so EaSe has advantages on time and token cost (see Ours in LABEL:fig:banner).
We evaluate EaSe on the widely used ScanRefer (Chen et al., 2020) and Nr3D (Achlioptas et al., 2020) datasets. Experimental results show that EaSe  achieves 52.9% accuracy on Nr3D and 49.2% Acc@0.25 on ScanRefer, matching the performance of agent-based approaches (Xu et al., ; Fang et al., 2024) while offering significant advantages in grounding time and token cost. In paticular, on Nr3D, EaSe  outperforms VLMGrounder (Xu et al., ) in accuracy while requiring less than $1/20$ of the grounding time and less than $1/6$ of the token usage. In addition, EaSe  significantly outperforms ZSVG3D (Yuan et al., 2024b) in accuracy with comparable grounding time and token cost. In conclusion, among various training-free 3DVG methods, EaSe  strikes an excellent balance between accuracy and efficiency.
Figure 1: Overview of EaSe. Block (a) and (b) compare the difference of relation encoders between EaSe and previous neuro-symbolic approaches. Relation encoders compute relation features using object positions from the scene. The executor executes the symbolic expression with relation features and gets candidate objects. Block (c) demonstrates a VLM selects the target from candidate objects given scan images containing them.
2 2. Related Works
3D Visual Grounding
3D visual grounding focuses on finding an object in 3D scene based on descriptions to appearance or location. ScanRefer (Chen et al., 2020) and ReferIt3D (Achlioptas et al., 2020) are two popular benchmarks on this task, providing rich and diverse object-utterance pairs on ScanNet (Dai et al., 2017). Many supervised approaches train a end-to-end model for 3DVG. Jain et al. (2022) combines bottom-up and top-down object detection and grounding with transformer. Chen et al. (2022) designs fine-grained neural spatial relation encoders. Ziyu et al. (2023) achieves promising performance by training on large scale well annotated dataset. Compared to supervised approaches, EaSe doesn’t rely on large scale training data. Neuro-symbolic approaches (Hsu et al., 2023; Feng et al., 2024; Li et al., 2024b) use LLMs to parse the referring utterance into symbolic expressions and learnable neural networks as spatial relation encoders. In contrast, EaSe use Python codes are spatial relation encoders.
Yuan et al. (2024b) manually annotates functions and prompt LLM to generate function calling code. Yang et al. (2024); Fang et al. (2024) create agents where LLM or VLM can reason and take actions iteratively. Xu et al. uses VLM and images from the scene to figure out the target object without the need for detection or segmentation modules. Inference cost of EaSe is as low as Yuan et al. (2024b) and the performance is comparable as Yang et al. (2024); Xu et al. ; Fang et al. (2024).
LLM Programming
LLM can generate executable codes (Roziere et al., 2023; Zhu et al., 2024) from scratch or according to in context documents.  Li et al. (2023a) executes generated codes for precise mathematical reasoning.  Liang et al. (2023) uses LLM to generate policy codes for robot controlling.  Xie et al. ; Ma et al. (2023) prompt LLM to design reward functions for RL training and use feedback from real world to improve them. Zhou et al. (2024) generates functions to refine pre-training corpora. Gupta and Kembhavi (2023); Yuan et al. (2024b) annotate tool APIs and use LLM to convert natural language instructions into API-calling codes. LLMs are also able to refine its generated code based on error message from executor or feedback from environment. Le et al. (2022); Chen et al. (2023) use the feedback from the outside environment to fine-tuning language models, Ma et al. (2023) uses RL training trajectories to select and improve its reward functions. In 3DVG area, Yuan et al. (2024b); Fang et al. (2024) also use code to process spatial relations, But to the best of our knowledge, we are the first to design specific test suites for code selection and improvement.
Neuro-symbolic Reasoning
Neuro-symbolic reasoning methods parse natural language to symbolic expression and get reasoning results by executing it with neural networks. Li et al. (2023b) provides a neuro-symbolic interface that supports many logical reasoning and seamless integration with other pre-trained model seamlessly. Cheng et al. (2023) uses LLM query to assist SQL execution. Mao et al. (2019); Hsu et al. (2023); Feng et al. (2024) define domain-specific language and use neural networks as relation encoders for 3DVG. EaSe uses similar symbolic frameworks, but our relation encoders are Python codes. Spatial knowledge of our encoders is from LLM instead of 3D training data.
3 3. Method
3.1 Problem Statement
3D visual grounding tasks involve a scene, denoted as $\mathcal{S}$, represented by an RGB-colored point cloud containing $C$ points, where $\mathcal{S}\in\mathbb{R}^{C\times 6}$. Accompanying this is an utterance $\mathcal{U}$ that describes an object within the scene $\mathcal{S}$. The objective is to identify the location of the target object $\mathcal{T}$ in the form of a 3D bounding box. In the ReferIt3D dataset (Achlioptas et al., 2020), bounding boxes for all objects are provided, making the visual grounding process a task of matching these bounding boxes to the scene $\mathcal{S}$. In contrast, the ScanRefer dataset (Chen et al., 2020) provides only the scene point cloud, requiring additional detection or segmentation modules to accomplish the grounding task.
3.2 Grounding Pipeline
We adhere to the previous SOTA neuro-symbolic framework for 3DVG (Hsu et al., 2023; Feng et al., 2024). A semantic parser (LLM) converts $\mathcal{U}$ into a symbolic expression $\mathcal{E}$ in JSON. Relation encoders compute the features for relations such as under and right within $\mathcal{E}$. Finally, the features, along with category features of relevant object classes in $\mathcal{E}$ (“chair" and “desk"). These features and logits are subsequently used to calculate the matching degrees between $\mathcal{S}$ and the objects based on $\mathcal{E}$.
Semantic Parsing
We employ GPT-4o (OpenAI, 2024) to parse $\mathcal{U}$ into JSON expression $\mathcal{E}$, which contains the categories and spatial relations in $\mathcal{U}$. For example, the utterance “chair near the table" can be represented as:
{"category": "chair", "relations":
[{"relation_name": "near",
"objects": [{"category": "table"}]}]}
Human-annotated natural language expressions exhibit diverse descriptions of relations, leading to a long-tail distribution of relation_name in parsed expressions. To mitigate this, we predefine a set of common relation names and prompt LLM to select from them for $\mathcal{E}$ instead of using the original terms from $\mathcal{U}$.
Based on the number of associated objects, the relations can be categorized into unary, binary, and ternary (Feng et al., 2024). For simplicity, attributes that describe properties of a single object, such as “large" or “at the corner" are treated as special types of unary relations. We presents our predefined set of relations along with their classifications in Table 1.
Table 1: Classification of all relations.
Classification Relations
unary large, small, high, low, on the floor,
against the wall, at the corner
binary near, far, above, below,
left, right, front, behind
ternary between
Feature Computing
Neuro-symbolic approaches(Hsu et al., 2023; Feng et al., 2024) uses neural network as relation encoders. In contrast, our relation encoders are Python codes generated by LLMs. It take the scene’s point cloud data and object segmentation as input and compute spatial relation features of objects in an explicit way. The object category features are from a pre-trained point cloud classifier. The feature of unary relations and category features, denoted as $f_{\text{unary}}\in\mathbb{R}^{N}$, can be seen as the similarity between objects and their respective categories or relations, where $N$ is the number of objects in the scene. Features of binary relations $f_{\text{binary}}\in\mathbb{R}^{N\times N}$ represent the likelihood of binary relations existing between all possible pairs of objects. For instance, the element $f_{\text{near}}^{i,j}$ quantifies the probability that the $i$-th object is “near" the $j$-th object. The structure of ternary features follows a similar pattern, capturing relations involving three objects.
Execution
Our executor is similar to Hsu et al. (2023). The executor utilizes the symbolic expression $\mathcal{E}$ and the associated relation features to identify the target object $\mathcal{T}$. Since the elements in the features represent probabilities or corresponding relations/categories, logical conjunctions within $\mathcal{E}$ are modeled using product operations.
For a given symbolic expression $\mathcal{E}$: The executor use $\mathcal{E}$ and associated features to identify $\mathcal{T}$. Since elements in the features represent the probabilities or corresponding relation or category, the logical conjunction in $\mathcal{E}$ can be represented through the product operation. For a symbolic expression $\mathcal{E}$, its category feature $f_{\text{category}}$ of calculated initially by a classifier. Subsequently, the executor processes each relation individually of the relations field. For the $i$-th relation, the relation feature $f_{\text{relation}}$ is computed and multiplied with $f_{\text{category}}$ of its related objects, resulting a $f_{\text{i}}\in\mathbb{R}^{N}$. After processing all relations, we obtain multiple relation features, each of size $\mathbb{R}^{N}$. They are aggregated with $f_{\text{category}}$ through an element-wise product to compute the final matching degree.
3.3 Relation Encoders
The sizes and positions of objects in 3D scenes are intrinsically linked to specific spatial relations. For instance, “near” depends on the distance between objects, while “large” depends on their volume. In EaSe , each of these spatial relations is handled by a dedicated piece of Python code that computes its associated feature.
The generation and optimization of these codes are illustrated in Figure 2. During the initial code generation stage, an optimized snippet for the relevant relation is retrieved as an in-context example (§$\S$ 3.3.2). Then we sample multiple responses from LLM and test them with a test suite (§$\S$ 3.3.3). Any failures in the test suite trigger error message synthesis, guiding the LLM to iteratively refine and optimize the codes (§$\S$ 3.3.4). The following sections provide a more detailed explanation of this process.
3.3.1 Class Schema
For each relation, we define a corresponding Python class as its encoder. The class is initialized with the scene’s point cloud data and object segmentation and provides two key methods: _init_param, which computes the necessary parameters for feature derivation. For instance, in the “near” class, it calculates distances between each pair of objects; forward: which performs numerical operations such as inversion or exponentiation and returns the relation feature.
3.3.2 In Context Example
Adding in-context example (ICE) into prompt can significantly improve the accuracy of responses from LLM (Brown et al., 2020). To reduce human effort and provide suitable ICE for different relation encoders’ generation, we utilize a semantic based retrieval strategy. For example, relation encoders for “near" and “far" may both compute pairwise distances but differ only in the numerical processing, so the well optimized relation encoder for “near" can serve as ICE for generation of “far".
We construct a directed acyclic graph $G$ (see Figure 6). Each node in $G$ represents a specific relation, and an edge from node A to node B indicates that the encoder for relation A is used as an ICE when generating for relation B. With this structure, the entire code generation process follows a topological ordering over $G$ automatically.
3.3.3 Unit Test
LLMs may not always generate a perfect code within one attempt (Olausson et al., 2023). Wu et al. (2024) proposes that more sampling increases the likelihood of correct answers. So we design unit test suites to select the plausible relation encoders from multiple LLM responses.
Take binary relation “above" for example, we collect a few triplets in the format of target object, same class distractor, anchor object from training set, with each triplet serving as a unit test case. The scale for each relation is small (less than 50 for most of relations). In the relation feature from encoder, if the distractor-anchor pair (probability of distractor “above" anchor) is larger than the target-anchor pair, the test fails and an error message like [target bbox] is above [anchor bbox] So metric value of [target bbox] "above" [anchor bbox] should be larger than than the metric value of [distractor bbox] "above" [anchor bbox]. is synthesized. 3 shows an example of the error message.
3.3.4 Code Generation and Optimization
As illusrated in Figure 2, for any relation, we begin by prompting the LLM with a high-level task description, the relation name, and an in-context example retrieved from codes that are already optimized as §$\S$ 3.3.2. Then we sample $N_{\text{sample}}$ candidate code snippets from LLM, where $N_{\text{sample}}$ is a configurable hyperparameter.
Next, each generated code is tested using the unit tests defined in §$\S$ 3.3.2. We select the $top_{k}$ codes that pass the most test cases and subject them to an optimization phase. During optimization phase, the LLM is given the initial prompt, the code to be optimized, and error message produced by the test suite. It then revises the code according to these errors.
This procedure of testing and refinement is repeated for up to $N_{\text{iter}}$ iterations. Ultimately, we adopt the code that achieves the highest pass rate across all test cases. In the event of a tie, we select the code that underwent more optimization steps. The optimization and selection algorithm is shown in Algorithm 1.
1
Require : relation name $R$, relation name $G$, code library $L$, test cases $C$, LLM LLM, test suites $T$, initial prompt prompt
Output : best_code
2
Hyperparameters : search iteration $N$, sample number $M$, evolving example number $K$
3
1
[…]
5for $j\leftarrow 1\dots M$ do
       $acc_{j},err_{j}\leftarrow T(F_{j})$
        // Test each code.
6
7
[…]
11for $i\leftarrow 2\dots N$ do
12
13      $\texttt{results}\leftarrow[]$
14      for $j\leftarrow 1\dots K$ do
15             $(F_{\text{old}},err_{\text{old}})\leftarrow\texttt{TopK}[j]$
16            $\texttt{prompt}_{\text{ref}}\leftarrow\texttt{init\_prompt}+F_{\text{old}}+err_{\text{old}}$
17            $F_{1},\dots,F_{M}\leftarrow\texttt{LLM}(R,\texttt{prompt}_{\text{ref}})$
18            for $k\leftarrow 1\dots M$ do
19                   $\texttt{results}.\texttt{append}(F_{k})$
20
21
[…]
23      foreach $F_{k}\in\texttt{results}$ do
24             $acc_{k},err_{k}\leftarrow T(F_{k})$
25            if $acc_{k}=1$ then
26                   return $F_{k}$
27
28            if $_{k}>\texttt{max\_acc}$_acc then
[…]
31
[…]
33
[…]
35
[…]
return best_code
Algorithm 1 Code Generation and Optimization
3.4 VLM Decision
The visual information like description on color or shape in utterances can be essential for accurate grounding, particularly for natural datasets like Nr3D (Achlioptas et al., 2020) and ScanRefer (Chen et al., 2020). When two candidate objects share similar class and spatial position, visual perception is required to distinguish between them.
Following Xu et al. , we incorporate GPT-4o to identify the target object from a set of candidates by utilizing 2D images from ScanNet (Dai et al., 2017) as additional context. Specifically, we select the top five objects based on output from executor and retain those whose logits exceed a chosen threshold as final candidates. Next, we retrieve the corresponding images from ScanNet (Dai et al., 2017) by matching each candidate’s point cloud to camera parameters, thereby finding images that include the candidate objects. Out of these, we select eight images that contain the most prominent views of each candidate (i.e., the largest bounding boxes in the image). We then annotate each image with object IDs and stitch them together in a $4\times 2$ grid. Finally, we prompt GPT-4o to identify the target object from stitched images. By integrating these visual cues, the VLM decision step effectively disambiguates candidates that appear similar in terms of class and spatial attributes, yielding more accurate results.
Figure 2: Overview of the generation and optimization process of relation encoders.
4 4. Experiments
4.1 Experimental Settings
Dataset
We conduct experiments on Nr3D subset from ReferIt3D (Achlioptas et al., 2020) dataset and ScanRefer (Chen et al., 2020). ReferIt3D has 2 subsets: Nr3D and Sr3D. Nr3D subset utterances contain human-annotated utterances and Sr3D contains synthesized ones. Based on the number of same-class distractors, the dataset can be categorized into “easy" and “hard" subsets depending on the number of same class distractors. The easy subset has a single distractor and the hard subset has multiple. The dataset can also be split into “view dependent" and “view independent" subsets according to if there are some keywords in the utterance. Ground truth object bounding boxes are given in ReferIt3D default test setting. So the metric is an exact match between the predicted bounding box and the target bounding box. In ScanRefer, no GT object mask is provided, the evaluation metric is the intersection over union (IoU) value between predicted bounding box and GT bounding box. There are two different thresholds: 0.25 and 0.5, so the metric names are Acc@0.25, Acc@0.5 respectively.
Implementation Details
For code evolving, we set $N_{sample}$ and $N_{iter}$ to 5, $topk$ to 3. We mainly use gpt-4o-2024-08-06 model with a temperature of 1.0 and top_p of 0.95. For a fair comparison, we use the same object classification result and evaluation code as  Yuan et al. (2024b) for evaluation of ReferIt3D. For evaluation of ScanRefer, we use the object detection and classification results of (Jain et al., 2024). (internal experiments show that EaSe has similar performance when using same object proposals as Yuan et al. (2024b).) For VLM decision, we use the same temperature and top_p as Xu et al. , the threshold is 0.9 for Nr3D, and 0.1 for ScanRefer.
Baselines
We compare EaSe against both supervised and training-free methods, evaluating their performance and inference cost. The supervised baselines include BUTD-DETR (Jain et al., 2022), Vil3DRel (Chen et al., 2022), 3D-VisTA (Ziyu et al., 2023), and CoT3DRef (Bakr et al., 2023), while the training-free approaches include ZSVG3D (Yuan et al., 2024b), VLM-Grounder (Xu et al., ), and Transcrib3D (Fang et al., 2024).
For the Nr3D dataset, Transcrib3D (Fang et al., 2024) uses ground-truth object labels, which are substantially more accurate than the predicted labels we employ. Additionally, the neuro-symbolic method NS3D (Hsu et al., 2023) operates under a more limited evaluation protocol; therefore, we compare to them within their specific settings.
4.2 Quantitative Results
ReferIt3D
Table 4 presents the results on Nr3D. Compared to other training-free baselines, EaSe achieves better overall accuracy than both ZSVG3D (Yuan et al., 2024b) and VLM-Grounder (Xu et al., ). When combined with VLM, EaSe further narrows the gap in overall performance relative to the supervised BUTD-DETR (Jain et al., 2022), especially on the view-dependent (VD) subset. However, it still lags behind other more recent supervised methods, which are trained on large scaled specific 3D datasets. We evaluate EaSe in the settings of Fang et al. (2024) and Hsu et al. (2023) respectively. In Fang et al. (2024)’s setting, the GT object labels are utilized for more accurate category-level object recognition. EaSe underperforms Fang et al. (2024) by 2.4% (67.8% vs. 70.2%). The high accuracy of Fang et al. (2024) comes from LLM based reasoning during inference time, which greatly hurts its inference efficiency. In the setting of  Hsu et al. (2023), EaSe outperforms it by 7.4% without any training.
ScanRefer
Table 3 shows the results on ScanRefer. VLM-Grounder (Xu et al., ) is evaluated on a randomly selected subset containing 250 samples EaSe outperforms ZSVG3D (Yuan et al., 2024b) by $12.8\%$ (49.2% vs. 36.4%) when leveraging a VLM. Moreover, the performance gap between EaSe (49.2%) and Xu et al. (51.6%), Fang et al. (2024) (51.3%) is relatively small, demonstrating that EaSe balances efficiency with competitive accuracy.
Grounding Cost
Table 2 compares the inference time and token consumption of various training-free methods on Nr3D. We randomly sampled a subset of the Nr3D test set to calculate the average inference time and token usage. Agent-based methods (Xu et al., ; Fang et al., 2024) rely on multiple calls to the LLM during a single inference, with the context continuously growing. As a result, these methods consume significantly more time and tokens (27.0s and 50.3s, respectively) compared to EaSe (2.1s).
In contrast, methods like EaSe (without VLM) and ZSVG3D (Yuan et al., 2024b) only require a single call to the LLM for semantic parsing with a short prompt, leading to lower time and token consumption during inference, and EaSe significantly outperforms ZSVG3D in accuracy by a large margin. When combined with a VLM, EaSe still achieves more than three times faster inference and significantly reduced token consumption compared to agent-based methods (Xu et al., ; Fang et al., 2024).
In conclusion, the above quantitative results highlight that EaSe effectively balances performance and efficiency, achieving leading accuracy alongside exceptional computational cost savings among training-free methods.
Table 2: Grounding time and token cost on Nr3D for training-free method that utilized LLM. EaSe has significant advantage, especially when compared to agent based methods (VLM-Grounder and Transcrib3D). $\dagger$: evaluated on a subset having 250 samples.
Method Time/s Token
ZSVG3D 2.4 2.5k
[…]
Transcrib3D 27.0 12k
EaSe (w/o VLM) 2.1 1.2k
EaSe 7.7 (+5.6) 3.1k (+1.9k)
Table 3: Performance on ScanRefer. EaSe has close performance with VLM-Grounder and BUTD-DETR. $\dagger$: evaluated on a subset having 250 samples.
Method Acc@0.25
BUTD-DETR 52.2
[…]
Transcrib3D 51.3
EaSe 49.2
ZSVG3D 36.4
Table 4: Performance on Nr3D. VD and VID stand for view-dependent and view-independent, respectively. We only report time and token consuming for training-free methods which involve LLM or VLM. EaSe outperforms other training-free baselines in overall performance, and has significant advantages in terms of time and token consumptions. For comparison with Transcrib3D, we use GT object labels. For comparison with NS3D, we evaluate EaSe on the same Nr3D subset as NS3D. $\dagger$: VLM-Grounder is evaluated on a subset having 250 samples. * : we re-run ZSVG3D using GPT-4o.
Method Overall Easy Hard VD VID
ViL3DRel 64.4 70.2 57.4 62.0 64.5
CoT3DRef 64.4 70.0 59.2 61.9 65.7
3D-VisTA 64.2 72.1 56.7 61.5 65.1
BUTD-DETR 54.6 60.7 48.4 46.0 58.0
ZSVG3D 40.2 49.1 31.1 37.8 41.6
[…]
EaSe (w/o VLM) 50.7 58.7 43.0 45.6 53.2
EaSe 52.9 - - - -
Transcrib3D 70.2 79.7 60.3 60.1 75.4
EaSe (w/o VLM) 65.7 75.6 56.2 58.7 69.1
EaSe 67.8 - - - -
NS3D 52.8 - - - -
EaSe (w/o VLM) 60.2 - - - -
4.3 Qualitative Results
Scene Visualization
Figure 3: Explanation and visualization of the grounding steps. Anchors (desk for the upper case and door for the lower) are marked by red circle. We only visualize a part of objects that match the below conditions well in green, objects of brighter color have higher scores of meeting the condition.
In Figure 3, we present two step-by-step grounding process of EaSe , illustrating how the final grounding results are constructed through combinition of multiple conditions within the referring utterance. The corresponding utterance for the second row is, “When facing the door, it’s the shelf above the desk on the right." The execution of the symbolic expression involves four steps, progressing from left to right. Objects with higher scores are highlighted in green, where lighter green indicates closer match between an object the condition below. First, we can find objects “right of the door" by the category feature of “door" and relation feature “right", which are in green color. Then we can find “desk right of the door" from them using the category feature of “desk". After that, objects which satisfied “above desk and on door’s right" can be found by joint condition of “above desk" and “right of the door". Finally the target “shelf" is grounded from them. More visualization results can be found in the Figure 8.
4.4 Ablation Studies
Figure 4: The learning curves of different variants. The x-axis is the generation number of the program. The y-axis is the accuracy of the program on the test set. It is worth noting that the curves of our full model Refer to caption coincide with the curves of Variant 3 Refer to caption in the second row (“Left”, “Above”, and “Corner”) since Variant 3 can be regarded as the first stage of generating all programs.
We conduct ablation studies to investigate the impact of various components during the code generation process, evaluating three different variants. We chose to analyze six relations that are relatively complex and have multiple iterations during the generation and optimization process. (For relatively simple relations like “small", the generated code passes all unit tests in the first generation so no optimization is used.) In all three variants, no in-context example (§$\S$3.3.2) is provided. The three different variants are: 1) Generation only: In this variant, we generate the code by directly prompting the LLM and sample multiple codes. The code with the highest pass rate on unit tests is selected. 2) Self-optimization only: In this variant, we use the same prompt as variant 1) and replace the feedback message (§$\S$3.3.4) containing specific failure cases with a general optimization instruction. 3) No ICE: This variant has feedback message §$\S$3.3.4 and no in-context example is used for generation. For the relations that no in-context examples available(“left”, “above”, and “corner”), variant 3 is identical to the first stage of full version of EaSe , so we only plot variants 1 and 2 on the corresponding plots. For relations “right", “below" and “between", we conduct experiment on all 3 variants. To control for the impact of the initial generation, we use the same results of iteration 0 across variant 2 and variant 3.
Figure 4 illustrates the results of ablation study, different variants are represented by lines of different colors. The horizontal axis represents the number of iterations, after an iteration ends, we choose a code based on their unit test pass rates and then evaluate it on the test set. The vertical axis shows the number of correctly solved test examples associated with the relation. We normalize them by dividing them by the maximum value. We can see that random sampling (Variant 1) shows comparable performance to the full model only for the “corner" and “between" relations. In Variant 2, which uses simple self-reflection, there is noticeable performance improvement across most relations except "corner." This variant achieves performance on "above" and "corner" close to that of the full model, as these relations are easier to generate. However, it struggles with the more complex relations. Variant 3, which incorporates feedback messages, improves upon the performance of Variant 2. For all relations except “right", it achieves results similar to the full model. However, for “right," “between," and “below" — where in-context examples are used to assist generation — Variant 3 shows a significant performance gap in the early iterations compared to the full model, highlighting the impact of in-context examples.
5 5. Conclusion
In this work, we propose EaSe , a training free method that uses Python code as relation encoder for 3DVG and an automatic pipeline for code generation and optimization. EaSe does not rely on large-scale data and has promising advantages in performance, and inference cost compared to other similar methods. We directly use knowledge LLM for relation encoder instead of human annotation or supervised learning.
However, some limitations still remain. In referring utterances, there are diverse referring utterances containing words other than objects and relations. But our method mainly focuses on objects and spatial relations so it cannot handle the non-relation words in some complex cases. Besides, we treat ignore the objects’ orientations, and global information of the scene.
[…]
Appendix A Prompts
In this section, we show the prompts we used.
Listing 1: Prompt for semantic parsing
You are a skilled assistant with expertise in semantic parsing.
## Task Overview
I will provide you with a sentence that describes the location of an object within a scene. Your task is to convert this description into a JSON format that captures the essential details of the object.
### The JSON object should include:
- **"category"**: string, representing the objects category.
- **"relations"**: a list of relationships between the object and other elements in the scene. Each relationship should be represented as a dictionary with the following fields:
- **"relation_name"**: string, specifying the type of relationship. The relationship can be:
- *Unary*: choose from [’corner’, on the floor’, against wall’, smaller’, larger’, taller’, lower’, within’].
- *Binary*: choose from [’above’, below’, beside’, close’, far’, left’, right’, front’, behind’, across’].
- *Ternary*: choose from [’between’, center’, middle’].
Only consider **simple** and **general** relations, donot make complex ones like "left of a blue box", "with dark appearance", "facing the window", etc. You should handle these by logical structures.
If the relationship is not mentioned in the list, you should choose the most appropriate relation above. **Never** create a new relation name!
- **"objects"**: a list of objects involved in the relationship. Every object in the list should have the same JSON structure. The list structure depends on the relationship type:
- *Unary*: The list should be empty.
- *Binary*: The list should contain one object.
- *Ternary*: The list should contain two objects.
- **"negative"**: boolean, indicating if the object is explicitly described as not having this relationship. Set this to True if applicable.
## Guidelines:
- First, generate a plan outlining the objects appearance and relationships based on the sentence. Then, use this plan to create the JSON representation.
## Examples:
### Example 1:
**Sentence**: The correct whiteboard is the one on a table.
**Plan**: "Correct" does not describe appearance. The appearances are "whiteboard" and "table", and the "whiteboard" is on the "table".
**Parsed JSON**:
“‘json
"category": "whiteboard",
"relations": [
"relation_name": "above",
"objects": [
"category": "table",
"relations": []
2 more examples.
Listing 2: An example of initial prompt for relation encoder generation
You are an expert on spatial relation analysis and code generation.
# Introduction to task
Your task is to write a Python class which can be used to compute the metric value of the existence of some spatial relationship between two objects in a 3D scene given their positions and sizes. Higher the metric value, higher the probability of the two objects have that relation.
In the class, you will be given the positions and sizes of the objects in the scene. The class should have a method forward which returns a tensor of shape (N, N), where element (i, j) is the metric value of the relation between object i and object j.
In the 3D scene, x-y plane is the horizontal plane, z-axis is the vertical axis.
# Introduction to programming environment
Here is an example class for Left relation. The class you write should have the same structure as the example class.
“‘python
class Left:
Make sure all tensors are placed on DEVICE‘, which has been defined in the environment.
The code output should be formatted as a python code string: "“‘python “‘".
# Some helpful tips
(1) You should only use the given variables, and you should not introduce new variables.
(2) The metric value should be sensitive to the input arguments, which means if the arguments change a little, the value should change a lot.
(3) The metric value should be 0 if the two objects dont have that relation, never set negative values!
(4) Never treat an object as its center point, you must consider the size of the bounding box, just like the example code. Never set an threshold to determine the relation. The value of the relation should be continuous and sparse.
(5) You should imagine that you are at position (0, 0) to determine the relative positions.
(6) Remember you are **in** the scene and look around, not look from the top. So never use the same way as 2D environment.
Propose your method first and then generate the code. Think step by step.
Dont use any axis or specific direction as the reference direction or right direction, your method should work for any perspectives.
Listing 3: An example for feedback message
We have run your code on some cases. Here are 3 failure cases:
# Case 1.
Metric value of object tensor([ 0.3992, -0.5619, 0.8831, 0.3921, 0.3476, 0.1059], device=’mps:0’) "above" object tensor([-0.0432, -0.6965, 0.8483, 0.6526, 0.4943, 0.3061], device=’mps:0’) should be larger than 0. Metric value of object tensor([ 0.3992, -0.5619, 0.8831, 0.3921, 0.3476, 0.1059], device=’mps:0’) "above" object tensor([-0.0432, -0.6965, 0.8483, 0.6526, 0.4943, 0.3061], device=’mps:0’) should be higher than the metric value of object tensor([0.5338, 1.1607, 1.1160, 0.2121, 0.3323, 0.8192], device=’mps:0’) "above" object tensor([-0.0432, -0.6965, 0.8483, 0.6526, 0.4943, 0.3061], device=’mps:0’).
more 2 cases
The first three are the center of the object, the last three are the size of the object. x-y is the horizontal plane and z is the vertical axis.
After test, the pass rate of your code is too low. So you MUST check carefully where the problem is. If you cant find the problem, you should
come up with a new algorithm and re-write your code.
Dont forget the following tips:
(1) You should imagine that you are at position (0, 0, 0) to determine the relative positions.
(2) Remember you are **in** the scene and look around, not look from the top. So never use the same way as 2D environment.
(3) Dont use any of x-axis or y-axis as your perspective, Your method should work for every perspective.
(4) The horizontal plane is x-y plane.
Please carefully analyze each of the failure case and explain why your code failed to pass it. The reason can be incorrect test case might or your code might not be able to handle some specific cases. Please write your analysis for each of the failure cases.
After the analysis of all cases, you should write the improved code based on your analysis. But **never** modify on the class methods and function parameters.
Some possible improvement ways:
1. Use a new algorithm to calculate the metric value rather than just modifying the existing code.
2. Consider carefully what other factors might be relevant to the spatial relationship between two objects and use them in your calculation.
3. Check the correctness of the input data and the calculation process.
Listing 4: The general prompt for self-reflection in variant 2 of ablation studies.
Reflect on the code above, think carefully how to make it better. For example, check if you ignore some factors that may affect the result or use a wrong method.
Then you must re-write the code in the same format. Remeber all the tips!
Appendix B Implementation Details
B.1 Definition of Symbolic Expression
The semantic parser translates the referring utterance $\mathcal{U}$ into a JSON-based symbolic expression $\mathcal{E}$ comprising the following elements:
  • Category: A string indicating the category of the target object referenced in $\mathcal{U}$.
  • Relations: A list specifying spatial constraints relative to the target object. Each entry in this list contains:
    • relation_name: A string identifying the spatial relation in $\mathcal{U}$ (e.g., “near,” “above”).
    • anchors: A list of objects that share the given spatial relation with the target object. Each object is represented as its own JSON entity.
    • negative: A boolean value which, if set to true, denotes that the target object should not exhibit the specified spatial relation.
Appendix C Definition of Symbolic Expression
C.1 Example of Relation Encoder
An example of relation encoder for relation “Near" is shown in Figure 5. The class is initialized using the bounding boxes of all objects in the scene. he function _init_params computes the distances between all object pairs. The forward function calculates the relation feature for "Near" by performing numerical operations on these distances.
class Near:
def __init__(
self,
object_locations: torch.Tensor) -> None:
Args:
object_locations: torch.Tensor, shape (N, 6),
N is the number of objects in the scene.
self.object_locations = object_locations.to(DEVICE)
self._init_params()
def _init_params(self) -> None:
Computing some necessary parameters about ‘Near‘ relation and
initialize ‘self.param‘.
# Based on average size to calculate a baseline distance
sizes = self.object_locations[:, 3:]
self.avg_size_norm = sizes.mean(dim=0).norm().to(DEVICE)
def forward(self) -> torch.Tensor:
Return a tensor of shape (N, N),
where element (i, j) is the metric value
of the ‘Near‘ relation between object
i and object j.
centers = self.object_locations[:, :3]
sizes = self.object_locations[:, 3:]
# Calculate the pairwise distances between object
centers in a vectorized manner
diff = centers.unsqueeze(1) - centers.unsqueeze(0)
distances = diff.norm(dim=2)
# Calculate the "nearness" metric based
on distances and average size norm
nearness_metric = torch.exp(-distances /
(self.avg_size_norm + 1e-6))
# Set diagonal to zero since an object
cannot be near itself
nearness_metric.fill_diagonal_(0)
return nearness_metric.to(DEVICE)
Figure 5: Example relation encoder for “Near".
C.1.1 In Context Example Selection
The graph used for select in context example for relation encoder generation is shown in Figure 6.
Figure 6: The graph for in context example selection. An edge from node A to node B means that the encoder for relation A is used as an in context example when generating for relation B.
Category Features
Yuan et al. (2024b) provides the predicted classified label for each object. For the category feature $f_{\text{category}}\in\mathbb{R}^{N}$ ($N$ is the number of objects), we compute the cosine similarity $sim\in\mathbb{R}^{N}$ between the category and the predicted labels using CLIP (Radford et al., 2021). Subsequently, we define the category feature as:
$f_{\text{category}}=softmax(100\times sim)$
To optimize computation, we cache features for different categories in a dictionary $C$.
Relation Features
We use the object bounding boxes in the scene to initialize the relation encoders and then call the forward() function to compute the corresponding relation feature, f_rel. These relation features are also cached in a dictionary $R$ for efficient reuse.
Execution Details
The detailed execution algorithm is presented in Algorithm 2, utilizing the precomputed category features and relation features. The Execute function runs recursively to compute the $\texttt{logits}\in\mathbb{R}^{N}$ ($N$ is the number of objects).
Appendix D More Quantitative Results
NS3D
We show evaluation results in NS3D(Hsu et al., 2023) in Table 4 . NS3D can only learn concepts (e.g. relation name, category name) from training set and its parsing results of Nr3D contains more than 5,000 concepts, resulting a long-tailed problem. So it selects a subset containing 1,041 test examples, which only contains the same concepts as Sr3D, the dataset it is mainly trained. On NS3D subset, EaSe have a large improvement than Ns3D (more than 7%), which shows the advantage of EaSe for processing natural grounding tasks.
Sr3D
In Table 5, we show the evaluation results on NS3D, a subset of Nr3D, proposed by Hsu et al. (2023). Even not using training data of Sr3D, EaSe still achieves comparable performance with Hsu et al. (2023) on both setting (w/ and w/o GT labels). We show evaluation result on Sr3D, a subset of Achlioptas et al. (2020) in Table 5. In the standard setting, If using GT object labels, accuracy of our method (w/o VLM) on Sr3D is 95.3%, performance of NS3D and Fang et al. (2024) are 96.9% and 98.4%. So we believe that the bottleneck of Sr3D performance is object detection and classification rather than spatial relation understanding because its relation annotation is synthesized by relatively simple functions. So we mainly focus on natural benchmarks (Nr3D and Scanrefer) which have complex and real spatial relations.
Table 5: Performance on Sr3D.
Method Overall NS3D
BUTD-DETR 67.0 -
NS3D (Hsu et al., 2023) 62.7 52.6
NS3D (Hsu et al., 2023) (w/ GT Label) 96.9 -
Transcrib3D (Fang et al., 2024) (w/ GT Label) 98.4 -
Ours (w/o VLM) 62.0 58.6
Ours (w/o VLM, w/ GT Label) 95.1 -
Appendix E More Qualitative Results
E.1 Effect of Code Optimization
We show the change between the initial response and the final code after multiple rounds of sampling and iterative refinement. By transitioning from a strict geometric overlap calculation to a continuous, exponential-based measure for both vertical and horizontal distances, the optimized code now captures nuances in “above” relationships more robustly. This improved formulation inherently handles scenarios where objects are close but not strictly overlapping, and it provides a more stable gradient for training. Consequently, the updated model passes all test cases by offering a smoother, more differentiable metric that better aligns with real-world spatial relations and passes more test cases.
class Above:
def __init__(self, object_locations: torch.Tensor) -> None:
Args:
object_locations: torch.Tensor, shape (N, 6),
N is the number of objects in the scene.
The first three columns are the center of the object (x, y, z),
the last three columns are the size of the object (width, height, depth).
self.object_locations = object_locations.to(DEVICE)
self._init_params()
def _init_params(self) -> None:
Compute necessary parameters about ‘OnTopOf‘
relation and initialize self.param if needed.
# shape (N, 3)
self.centers = self.object_locations[:, :3]
# shape (N, 3)
self.sizes = self.object_locations[:, 3:]
def forward(self) -> torch.Tensor:
Return a tensor of shape (N, N), where element (i, j) is the
metric value of the ‘OnTopOf‘ relation between object i and object j.
N = self.object_locations.shape[0]
# Expand dimensions for broadcasting
# Shape (N, 1, 3)
centers_i = self.centers.unsqueeze(1)
# Shape (N, 1, 3)
sizes_i = self.sizes.unsqueeze(1)
# Shape (1, N, 3)
centers_j = self.centers.unsqueeze(0)
# Shape (1, N, 3)
sizes_j = self.sizes.unsqueeze(0)
# Calculate vertical distances
vertical_distance = centers_i[..., 2] - centers_j[..., 2]
- (sizes_i[..., 1] + sizes_j[..., 1]) / 2
# Calculate horizontal overlaps
overlap_x = torch.clamp((sizes_i[..., 0] + sizes_j[..., 0]) / 2
- torch.abs(centers_i[..., 0] - centers_j[..., 0]), min=0)
overlap_y = torch.clamp((sizes_i[..., 2] + sizes_j[..., 2]) / 2
- torch.abs(centers_i[..., 1] - centers_j[..., 1]), min=0)
# Calculate the overlap area in the x-y plane
horizontal_overlap_area = overlap_x * overlap_y
# Calculate the metric value, only where object i is above object j
metric_value = torch.where(vertical_distance > 0,
horizontal_overlap_area / (1 + vertical_distance),
torch.tensor(0.0, device=DEVICE))
# Mask out the diagonal (i == j) to ensure
self-relations are not considered
metric_value.fill_diagonal_(0)
return metric_value
(a) Initial generation of relation encoder of “above".
class Above:
def __init__(
self,
object_locations: torch.Tensor) -> None:
Args:
object_locations: torch.Tensor, shape (N, 6),
N is the number of objects in the scene.
The first three columns are the center of the object (x, y, z),
the last three columns are the size of the object (width, height, depth).
self.object_locations = object_locations
self._init_params()
def _init_params(self) -> None:
Compute necessary parameters about ‘OnTopOf‘
relation and initialize parameters.
self.center = self.object_locations[:, :3] # (x, y, z)
self.size = self.object_locations[:, 3:] # (width, depth, height)
def forward(self) -> torch.Tensor:
Return a tensor of shape (N, N), where element (i, j)
is the metric value of the ‘OnTopOf‘
relation between object i and object j.
N = self.object_locations.shape[0]
# Calculate bottom of i and top of j
height_i = self.size[:, 2].view(N, 1)
bottom_of_i = self.center[:, 2].view(N, 1) - height_i / 2
top_of_j = self.center[:, 2].view(1, N)
+ self.size[:, 2].view(1, N) / 2
# Vertical proximity
vertical_proximity = torch.exp(-torch.abs(bottom_of_i - top_of_j)
/ (height_i / 2))
# Center distances
center_dist_x = torch.abs(self.center[:, 0].view(N, 1)
- self.center[:, 0].view(1, N))
center_dist_y = torch.abs(self.center[:, 1].view(N, 1)
- self.center[:, 1].view(1, N))
# Combined sizes
combined_size_x = (self.size[:, 0].view(N, 1) + self.size[:, 0].view(1, N)) / 2
combined_size_y = (self.size[:, 1].view(N, 1) + self.size[:, 1].view(1, N)) / 2
# Horizontal alignment
horizontal_alignment = torch.exp(-(center_dist_x / combined_size_x) - (center_dist_y / combined_size_y))
# Combine the metrics with a weight emphasizing more on vertical proximity
relation_metric = vertical_proximity * horizontal_alignment
# Remove self-relations by zeroing the diagonal
relation_metric.fill_diagonal_(0)
return relation_metric
(b) The relation encoder of “above" after multi-turn sampling and optimization
Figure 7: Example of code optimization result on relation encoder of “above". The left output is from the first iteration, which only pass 18 test cases out of 37. After 3 iterations of sampling and optimization, we get the the relation encoder as the right one. It passes all 37 test cases.
E.2 Relation Constraints
Feng et al. (2024) propose that certain spatial relations are symmetric, like “near" or “far", meaning that if object A is "near" B, then B should also be "near" A. Consequently, the features of these relations should be symmetric. Conversely, other relations are inherently asymmetric, such as “left" or “right". For these relations, if a feature element is positive (indicating the presence of the spatial relation), its corresponding symmetric element should be negative (indicating the absence of the reverse spatial relation).
Feng et al. (2024) leverage large language models (LLMs) to annotate these constraints and apply an auxiliary loss to enforce them during training. In contrast, while EaSe does not explicitly train or use specialized instructions to create these constraints, we observe that some LLM-generated relation encoders inherently produce relation features that satisfy these constraints. Moreover, for certain relations, these constraints are guaranteed due to the deterministic execution of the code. In Figure 10, we present four relation features for a scene:
  • Features for “near" and “far" are guaranteed to be symmetric.
  • For asymmetric features such as “left" and “right," if $f_{i,j}>0$, it is guaranteed that $f_{j,i}=0$.
E.3 Condition Level Accuracy
Our parsed symbolic expressions typically include one or more spatial conditions for the target object. However, some conditions in the referring utterance may be redundant.
For instance, if the referring utterance is "find the monitor on the floor and under the desk," and all monitors "on the floor" are also "under the desk," then one of these two conditions is redundant. This means that even if the method fails to process one of the conditions, it can still provide the correct grounding result.
To better understand EaSe ’s capability, we evaluate it on utterances containing a single condition.
We categorize objects of the same class into groups. Within each group, we collect the conditions for each object from the parsed expressions. Each condition is represented in JSON format, such as: "relation": ..., "anchors": [...].
These conditions are executed seamlessly to identify the best-matching object. We compute the average precision and recall for all condition-level matches. EaSe achieves an average precision of 67.5% and an average recall of 66.9%.
Require : symbolic expression $E$, category features $C$, relation features $R$
1
Output : logits
1
[…]
6
[…]
8
[…]
10
[…]
12
[…]
14
15
[…]
17             $\texttt{a}\leftarrow\texttt{Execute}(\texttt{anchors}[0])$
18
[…]
20
[…]
23
[…]
25
[…]
27
[…]
29
30
[…]
32
[…]
34
35            $\texttt{f}\leftarrow\texttt{max}(\texttt{f})-\texttt{f}$
36
[…]
38
39
Output : logits
Algorithm 2 Execute
E.4 More Visualization Results
We visualize more grounding examples in LABEL:fig:appendix. The first row illustrates the grounding process for the kitchen cabinet close to the fridge and beside the stove. In the process, the stove, objects beside the stove, and the objects near the fridge are sequentially grounded, culminating in the target kitchen cabinet highlighted in green. The second row shows the grounding process for right trash can below the sink. Starting with the objects below the sink, followed by the objects on the right of the sink, and finally combining these conditions to highlight the target trash can in green.
E.5 Effect of Unit tests
To demonstrate the impact of filtering generated code based on its accuracy on training cases, we selected six relations and plotted their performance. The x-axis represents the pass rate on training cases, while the y-axis shows the number of passed examples in all relevant test cases.
For straightforward relations such as “near" or “far", GPT-4o can pass all unit tests on the first attempt, so we focus on cases requiring multiple refinement steps.
The results, shown in Figure 9, indicate that for five out of six relations (excluding behind), the code with the highest pass rate on training cases achieves top-tier performance on the test set. However, for the behind relation, using the best-performing code on the training cases results in about 15 fewer passed test cases compared to using code with approximately 70% accuracy. Despite this, it still outperforms code with accuracy below 0.5.
This discrepancy for behind may stem from biases in the training data collection process. Overall, selecting code based on its performance on the training set is effective for achieving strong test set performance.
Figure 8: The target objects are: “Stove next to another stove and close to the fridge" (top row) and “Trashcan to the right of and below the sink" (bottom row).
Figure 9: Corresponding relation between the unit test pass rate and number of correct examples on test set.
Figure 10: Relation feature examples. The features of “near" and “far" are symmetric, meaning mutual relationships hold true in both directions. For “left" and “right," if an element is positive, its corresponding symmetric element is zero, ensuring asymmetry. Additionally, all diagonal elements are zero, as self-relations are not considered.