### v3

# Evolving Symbolic 3D Visual Grounder with Weakly Supervised Reflection

Abstract
3D visual grounding (3DVG) is challenging due to the need to understand 3D spatial relations. While supervised approaches have achieved superior performance, they are constrained by the scarcity and high cost of 3D vision-language datasets. Training-free approaches based on LLM/VLM agents eliminate the need for training data, but they incur prohibitive grounding time and token costs. To address the challenges, we introduce a novel training-free symbolic framework for 3D visual grounding, namely Evolvable Symbolic Visual Grounder (EaSe ). EaSe uses LLM-generated code to determine 3D spatial relations among objects and integrates VLMs to process their visual information. EaSe also implements an automatic pipeline that evaluates and optimizes the quality of generated code. Experimental results demonstrate that EaSe achieves 52.9% accuracy on Nr3D dataset and 49.2% Acc@0.25 on ScanRefer, ranking among the best training-free methods. Moreover, it substantially reduces the grounding time and token cost, offering a balanced trade-off between performance and efficiency. Code is available at https://github.com/OpenRobotLab/EaSe.
keywords:
3D Visual Grounding, Large Language Model, Code Generation
[…]
figures/teaser.pdf Comparision between two previous training-free 3DVG methods and our method(EaSe ). For a query, agent based methods employ multimodal LLM to process scene information. They are more accurate, but their online LLM generation increases time consumption. Visual programming (Visprog.) method uses offline annoated relation functions, thus reduces grounding time, but it doesn’t perform well. In contrast, EaSe utilizes offline LLM generation and optimization before grounding and improves relation functions to relation encoders. As a result, EaSe ’s accuracy is close to agents but it’s consumption is much lower.
1 1. Introduction
The 3D visual grounding (3DVG) task focuses on locating an object in a 3D scene based on a referring utterance. Numerous supervised methods have been proposed for 3DVG (Hsu et al., 2023; Jain et al., 2022; Huang et al., 2022; Chen et al., 2022; Huang et al., 2024; Ziyu et al., 2023; Bakr et al., 2023; Wu et al., 2023). These methods learn representations of referring utterances, object attributes and spatial relations from large scale 3D vision-language training data with high-quality annotations and achieve state-of-the-art performances on 3DVG. However, the scarcity of 3D vision-language datasets (Chen et al., 2020; Achlioptas et al., 2020), coupled with the high cost of their annotations, limits their applicability. Furthermore, some supervised methods are trained on these closed-vocabulary datasets, restricting their applicability in open-vocabulary scenarios.
In recent years, large language models (LLMs) and vision-language models (VLMs) have shown remarkable capabilities in reasoning, code generation, and visual perception. Building on these advancements, open-vocabulary and zero-shot 3DVG agents (Yang et al., 2023; Xu et al., ; Fang et al., 2024; Li et al., 2024a) are proposed. These methods let LLMs perform numerical computing and reasoning on object locations in the text modality (Yang et al., 2023; Fang et al., 2024), or let VLMs locate targets from scene scan images in the visual modality (Xu et al., ). Leveraging the reasoning and perceptual abilities of advanced LLMs and VLMs, these agents achieve superior accuracy compared to other training-free methods. However, they rely on LLMs to produce lengthy responses (containing planning, reasoning, or self-debugging processes) for every referring utterance. This online generation style results in significant costs in terms of grounding time and token usage (see Agents block in LABEL:fig:banner). In contrast, the visual programming method (Yuan et al., 2024b) utilizes LLM to generate a program which uses annotated relation functions and outputs the target object by executing the program. The generated program is short so its time and token consumption are much lower. However, it has trouble considering many spatial relations in the referring utterance simultaneously (Yuan et al., 2024a). This results in relatively low accuracy (see Visprog. in LABEL:fig:banner).
To address these dual challenges of accuracy and cost, we propose EaSe, a novel training-free symbolic framework that integrates LLMs and VLMs for 3D visual grounding, balancing both accuracy and inference cost. EaSe builds upon previous neuro-symbolic frameworks (Hsu et al., 2023; Feng et al., 2024) and uses Python code generated and optimized by LLMs as spatial relation encoders (see  Figure 1, block (b)). EaSe also employs a VLM to distinguish objects that differ only in visual appearance. Specifically, EaSe parses the referring utterance into a symbolic expression which encapsulates all mentioned object categories and their spatial relations. Given positions of all objects in the scene, the spatial relation encoders generate relation features which can represent spatial relations between them. Then an executor aggregates the symbolic expression, relation features and object categories to exclude most objects that do not match the referring utterance. Finally, the VLM identifies the target object from images of the remaining candidate objects (see Figure 1, block (c)). For more accurate spatial relation encoders, we generate them through iterative optimization processes instead of directly prompting the LLM. We introduce test suites to evaluate spatial relation encoders. The test suites not only enable us to select better relation encoders from LLM responses but also allow the LLM to leverage failed test cases to optimize its code. In contrast to the online generation of agent-based methods, our generation and optimization are performed offline, avoiding per-utterance code generation. The spatial relation encoders are reused in all grounding processes, so EaSe has advantages on time and token cost (see Ours in LABEL:fig:banner).
We evaluate EaSe on the widely used ScanRefer (Chen et al., 2020) and Nr3D (Achlioptas et al., 2020) datasets. Experiment results show that EaSe  achieves 52.9% accuracy on Nr3D and 49.2% Acc@0.25 on ScanRefer, matching the performance of agent-based approaches (Xu et al., ; Fang et al., 2024) while offering significant advantages in grounding time and token cost. On Nr3D, EaSe  outperforms VLMGrounder (Xu et al., ) in accuracy while requiring less than $1/20$ of the grounding time and less than $1/6$ of the token usage. In addition, EaSe  significantly outperforms ZSVG3D (Yuan et al., 2024b) in accuracy with comparable grounding time and token cost. In conclusion, among various training-free 3DVG methods, EaSe  strikes an excellent balance between accuracy and efficiency.
Figure 1: Overview of EaSe. Block (a) and (b) compare the difference of spatial relation encoders between EaSe and previous neuro-symbolic approaches. Spatial relation encoders compute relation features using object positions from the scene. The executor executes the symbolic expression with relation features and gets candidate objects. Block (c) demonstrates a VLM selects the target from candidate objects given scan images containing them.
2 2. Related Work
3D Visual Grounding
3D visual grounding (3DVG) aims to localize objects in 3D scenes based on natural language descriptions of appearance and spatial relations. Two dominant benchmarks, ScanRefer (Chen et al., 2020) and ReferIt3D (Achlioptas et al., 2020), leverage ScanNet (Dai et al., 2017) scenes to provide diverse object-utterance pairs. Supervised Methods typically train end-to-end models on annotated 3D vision-language data. Jain et al. (2022) integrates bottom-up object detection with transformer-based grounding, Chen et al. (2022) designs fine-grained neural networks to encode spatial relations. Though achieving promising accuracy, these methods suffer from expensive data annotation dependency (Ziyu et al., 2023). Neuro-Symbolic Methods (Hsu et al., 2023; Feng et al., 2024; Li et al., 2024b) attempt to mitigate data reliance by combining symbolic parsing with neural components. They parse referring utterances into symbolic expressions using LLMs and train neural networks as spatial relation encoders. Unlike these approaches, EaSe completely avoids training on large-scale 3D datasets. Training-free methods exploit pre-trained LLMs / VLMs for open-vocabulary 3DVG. Yuan et al. (2024b) uses LLMs to generate programs that call predefined functions to find the target object. Yang et al. (2024); Fang et al. (2024) deploy LLM/VLM-based agents that analyze object appearances and locations and find the target. Xu et al. uses VLMs and images from the scene to figure out the target object. Concurrently, Yuan et al. (2024a) proposes to replace the programming of Yuan et al. (2024b) by constraint satisfaction solving. Li et al. (2024a) parses landmark and perspective of the referring utterance and then uses VLMs to find the target object from a rendered image. Compared to these methods, EaSe offers a superior balance of accuracy and efficiency, outperforming  Yang et al. (2024); Xu et al. ; Fang et al. (2024) in grounding time and token cost while surpassing Yuan et al. (2024b) in accuracy.
LLM Programming
LLMs demonstrate growing proficiency in generating executable code (Roziere et al., 2023) for precise mathematical reasoning (Li et al., 2023), robotics control (Liang et al., 2023), tool use (Gupta and Kembhavi, 2023; Yuan et al., 2024b) or data cleaning (Zhou et al., 2024). Recent work further explores code refinement via environmental feedback, such as RL training trajectories (Ma et al., 2023) or real-world execution errors (Le et al., 2022; Chen et al., 2023). In the 3DVG area, Yuan et al. (2024b); Fang et al. (2024) also uses code to process spatial relations, but EaSe advances this paradigm by introducing test suites to automatically optimize code.
3 3. Method
3.1 Problem Statement
3D visual grounding tasks involve a scene, denoted as $\mathcal{S}$, represented by an RGB-colored point cloud containing $C$ points, where $\mathcal{S}\in\mathbb{R}^{C\times 6}$. Associated with this is an utterance $\mathcal{U}$ that describes an object within the scene $\mathcal{S}$. The objective is to identify the location of the target object $\mathcal{T}$ in the form of a 3D bounding box. In the ReferIt3D dataset (Achlioptas et al., 2020), bounding boxes for all objects are provided, making the visual grounding process a task of matching these bounding boxes to the scene $\mathcal{S}$. In contrast, the ScanRefer dataset (Chen et al., 2020) provides only the point cloud of the scene, requiring additional detection or segmentation modules to accomplish the grounding task.
3.2 Grounding Pipeline
We adhere to the previous SOTA neuro-symbolic framework for 3DVG (Hsu et al., 2023; Feng et al., 2024). A semantic parser (LLM) converts $\mathcal{U}$ into a symbolic expression $\mathcal{E}$ in JSON. Spatial relation encoders compute the relation features such as under and right within $\mathcal{E}$. Finally, the relation features, along with category features of object categories in $\mathcal{E}$, such as chair and desk are subsequently used to calculate the matching scores between $\mathcal{S}$ and the objects based on $\mathcal{E}$.
Semantic Parsing
We employ GPT-4o (OpenAI, 2024) to parse $\mathcal{U}$ into JSON expression $\mathcal{E}$, which contains the categories and spatial relations in $\mathcal{U}$. For example, the utterance “chair near the table” can be represented as:
{"category": "chair", "relations":
[{"relation_name": "near",
"objects": [{"category": "table"}]}]}
Human-annotated natural language expressions exhibit diverse descriptions of relations, leading to a long-tail distribution of relation_name in parsed expressions. To mitigate this, we define a set of common relation names and prompt LLM to select from them for $\mathcal{E}$ instead of using the original terms from $\mathcal{U}$.
Based on the number of associated objects, the relations can be categorized into unary, binary, and ternary (Feng et al., 2024). For simplicity, attributes that describe properties of a single object, such as “large” or “at the corner” are treated as special types of unary relations. We present our predefined set of relations along with their classifications in Table 1.
Table 1: Classification of all relations.
Classification Relations
unary large, small, high, low, on the floor,
against the wall, at the corner
binary near, far, above, below,
left, right, front, behind
ternary between
Feature Computing
Our spatial relation encoders are Python code generated by LLMs. They take the objects’ point clouds and positions as input and compute spatial relation features by explicit geometric calculations. The category features are from a pretrained point cloud classifier. Unary relation features $f_{\text{unary}}\in\mathbb{R}^{N}$ and category features $f_{\text{category}}\in\mathbb{R}^{N}$ measure the similarity between objects and their respective relations or categories, where $N$ is the number of objects in the scene. The features of the binary relation $f_{\text{binary}}\in\mathbb{R}^{N\times N}$ represent the likelihood that there are binary relations between all possible pairs of objects. For example, $f_{\text{near}}^{(i,j)}$ quantifies the probability that the $i$-th object is near the $j$-th object. Ternary features follow an analogous pattern for relations involving three objects.
Executor
Our executor follows a similar design to Hsu et al. (2023). Given the symbolic expression $\mathcal{E}$ and features, the executor computes the matching score between objects and referring utterance $\mathcal{U}$. Specifically, the category features $f_{\text{category}}$ are computed by an object classifier, and relation features are from our spatial relation encoders. For each relation in $\mathcal{E}$’s relations field, the corresponding relation feature $f_{\text{relation}}$ is multiplied with $f_{\text{category}}$ of its related objects, yielding intermediate features $\{f_{i}\in\mathbb{R}^{N}\}_{i=1}^{K}$ (where $K$ is the number of relations). Finally, all intermediate features and $f_{\text{category}}$ are aggregated via the element-wise product to compute the final matching score. See Algorithm 2 for more details.
3.3 Spatial Relation Encoders
Sizes and positions of objects in 3D scenes inherently determine spatial relations. For example, the near relation depends on pairwise distances, while large is determined by object volumes. In EaSe , each spatial relation is handled by a dedicated Python class (§$\S$ 3.3.1) that can compute its associated features given the object bounding boxes.
As illustrated in Figure 2, our encoder generation involves three phases: (1) Retrieving in-context examples from semantically similar relations (§$\S$ 3.3.2), (2) Sampling multiple code implementations from LLMs, and (3) Validating candidates through unit tests (§$\S$ 3.3.3). When test failures occur, we automatically synthesize error messages that guide iterative refinement (§$\S$ 3.3.4).
3.3.1 Class Schema
EaSe uses Python classes as spatial relation encoders. A class is initialized with the scene’s point cloud data and object segmentation and provides two key methods: _init_param, which computes the necessary parameters for feature derivation. For instance, in the “near” class, it calculates distances between each pair of objects; forward, which performs numerical operations such as inversion or exponentiation and returns the relation feature.
3.3.2 In Context Example
Adding in-context examples into the prompt can significantly improve the accuracy of responses from LLMs (Brown et al., 2020). To reduce human effort and provide suitable ICE for different relation encoders’ generation, we introduce a semantic-based retrieval strategy. For example, relation encoders for “near” and “far” may both compute pairwise distances but differ only in the numerical processing, so the well-optimized relation encoder for “near” can serve as an in-context example for the generation of “far”. We show the retrieval details in Figure 6.
3.3.3 Unit Test
Since LLMs may not always generate correct code in a single attempt (Olausson et al., 2023), inspired by Wu et al. (2024)’s finding that increased sampling enhances success probability, we design unit test suites to select the plausible relation encoders from multiple LLM responses.
Take the binary relation “above” as an example. We collect a few triplets in the format of target object, same class distractor, anchor object from the training set, with each triplet serving as a unit test case. The scale for each relation is small (less than 50 for most relations). In the generated relation feature for “above” $f$, if $f^{(distractor,anchor)}$ is larger than $f^{(target,anchor)}$, the test is deemed to have failed, and an error message like [target bbox] is above [anchor bbox] So metric value of [target bbox] "above" [anchor bbox] should be larger than the metric value of [distractor bbox] "above" [anchor bbox]. is synthesized. 3 provides an example of such an error message.
3.3.4 Code Generation and Optimization
As illustrated in Figure 2, for any relation, we begin by prompting the LLM with a high-level task description, the relation name, and an in-context example retrieved from codes that are already optimized as §$\S$ 3.3.2. Then we sample $N_{\text{sample}}$ candidate encoders from LLM, where $N_{\text{sample}}$ is a configurable hyperparameter. Next, each generated code is tested using the unit tests defined in § 3.3.3. We select the $top_{k}$ codes that pass the most test cases and subject them to an optimization phase. During the optimization phase, the LLM receives the initial prompt, the code to be optimized, and the error message produced by the test suite. It then revises the code according to these errors. This test and optimization procedure is repeated for up to $N_{\text{iter}}$ iterations. Ultimately, we adopt the code that achieves the highest pass rate across all test cases. In the event of a tie, we select the code that underwent more optimization steps. The optimization and selection algorithm is shown in Algorithm 1.
1
Require : relation name $R$, relation name $G$, code library $L$, test cases $C$, LLM LLM, test suites $T$, initial prompt prompt
Output : best_code
2
Hyperparameters : search iteration $N$, sample number $M$, optimizing example number $top_{k}$
3
1
[…]
5for $j\leftarrow 1\dots M$ do
       $acc_{j},err_{j}\leftarrow T(F_{j})$
        // Test each code.
6
7
[…]
11for $i\leftarrow 2\dots N$ do
12
13      $\texttt{results}\leftarrow[]$
14      for $j\leftarrow 1\dots K$ do
15             $(F_{\text{old}},err_{\text{old}})\leftarrow\texttt{TopK}[j]$
16            $\texttt{prompt}_{\text{ref}}\leftarrow\texttt{init\_prompt}+F_{\text{old}}+err_{\text{old}}$
17            $F_{1},\dots,F_{M}\leftarrow\texttt{LLM}(R,\texttt{prompt}_{\text{ref}})$
18            for $k\leftarrow 1\dots M$ do
19                   $\texttt{results}.\texttt{append}(F_{k})$
20
21
[…]
23      foreach $F_{k}\in\texttt{results}$ do
24             $acc_{k},err_{k}\leftarrow T(F_{k})$
25            if $acc_{k}=1$ then
26                   return $F_{k}$
27
28            if $_{k}>\texttt{max\_acc}$_acc then
[…]
31
[…]
33
[…]
35
[…]
return best_code
Algorithm 1 Code Generation and Optimization
3.4 VLM Decision
The visual information, like descriptions on color or shape in utterances, can be essential for accurate grounding, particularly for natural datasets like Nr3D (Achlioptas et al., 2020) and ScanRefer (Chen et al., 2020). When two candidate objects share a similar class and spatial position, visual perception is required to distinguish between them.
Following Xu et al. , we incorporate GPT-4o to identify the target object from a set of candidates by utilizing 2D images from ScanNet (Dai et al., 2017) as additional context.
Specifically, we select the top five objects based on the scores from the relation encoders and retain those whose logits exceed a chosen threshold as final candidates. Next, we retrieve the corresponding images from ScanNet (Dai et al., 2017) by matching each candidate’s point cloud to camera parameters, thereby finding images that include the candidate objects. Out of these, we select eight images that contain the largest projected bounding box area of each candidate. We then annotate each image with object IDs and stitch them together in a $4\times 2$ grid. Finally, we prompt GPT-4o to identify the target object from stitched images. By integrating these visual cues, the VLM decision step effectively disambiguates candidates that appear similar in terms of class and spatial attributes, yielding more accurate results.
Figure 2: Overview of the generation and optimization process of relation encoders.
4 4. Experiments
4.1 Experimental Settings
Dataset
We conduct experiments on the Nr3D subset of ReferIt3D (Achlioptas et al., 2020) dataset and ScanRefer (Chen et al., 2020). ReferIt3D has 2 subsets: Nr3D and Sr3D. The Nr3D subset utterances contain human-annotated utterances and the Sr3D contains synthesized ones. Based on the number of same-class distractors, the dataset can be categorized into “easy” and “hard” subsets. The easy subset has a single distractor, and the hard subset has multiple distractors. The dataset can also be split into “view dependent” and “view independent” subsets according to the referring utterance. Ground truth object bounding boxes are given in the ReferIt3D default evaluation setting. Therefore, the metric is an exact match between the predicted bounding box and the target bounding box. In ScanRefer, no GT object mask is provided; the evaluation metric is the intersection over union (IoU) value between the predicted bounding box and the GT bounding box. We use the Acc@0.25 metric here.
Implementation Details
For code optimization (§ 3.3.4), we set $N_{sample}$ and $N_{iter}$ to 5, $top_{k}$ to 3. We mainly use gpt-4o-2024-08-06 model with a temperature of 1.0 and top_p of 0.95. For a fair comparison, we use the object classification results from  Yuan et al. (2024b) for the evaluation of ReferIt3D. For evaluation of ScanRefer, we use the object detection and classification results of (Jain et al., 2024). (Experiments show that EaSe has similar accuracy when using object detection and classification results from Yuan et al. (2024b).) For VLM decision-making, we use the same temperature and top_p values as in Xu et al. . The threshold (§ 3.4 is set to 0.9 for Nr3D and 0.1 for ScanRefer.
Baselines
We compare EaSe against both supervised and training-free methods, evaluating accuracy, grounding time, and token cost. The supervised baselines include BUTD-DETR (Jain et al., 2022), Vil3DRel (Chen et al., 2022), 3D-VisTA (Ziyu et al., 2023), and CoT3DRef (Bakr et al., 2023), while the training-free approaches include ZSVG3D (Yuan et al., 2024b), VLM-Grounder (Xu et al., ), and Transcrib3D (Fang et al., 2024).
On the Nr3D dataset, Transcrib3D (Fang et al., 2024) uses ground-truth object labels, providing an advantage over methods which rely on predicted labels. Additionally, the neuro-symbolic method NS3D (Hsu et al., 2023) operates under a more limited evaluation protocol; therefore, we compare EaSe with them under their specific settings.
4.2 Quantitative Results
ReferIt3D
Table 4 presents the results on Nr3D. Compared to other training-free baselines, EaSe (without VLM) achieves higher overall accuracy than both ZSVG3D (Yuan et al., 2024b) and VLM-Grounder (Xu et al., ). When integrated with VLM, EaSe further narrows the gap in overall performance relative to the supervised BUTD-DETR (Jain et al., 2022), especially on the view-dependent (VD) subset. However, it still lags behind other more recent supervised methods, which are trained on large-scale task-specific 3D datasets. We further evaluate EaSe under the experimental settings of Fang et al. (2024) and Hsu et al. (2023) respectively. In Fang et al. (2024)’s setting, the ground truth (GT) object labels are utilized for more accurate category-level object recognition. EaSe slightly underperforms Fang et al. (2024) by 2.4% (67.8% vs. 70.2%), but EaSe has significant advantages in grounding time and token cost. Under the setting of  Hsu et al. (2023), EaSe outperforms it by 7.4% without requiring any training.
ScanRefer
Table 3 shows the results on ScanRefer. EaSe outperforms ZSVG3D (Yuan et al., 2024b) by $12.8\%$ (49.2% vs. 36.4%). Moreover, the performance gap between EaSe (49.2%) and Xu et al. (51.6%), Fang et al. (2024) (51.3%) is relatively small, demonstrating that EaSe balances cost with competitive accuracy.
Grounding Cost
Table 2 compares the average grounding time and token costs of various training-free methods on a randomly sampled subset of the Nr3D dataset. Agent-based methods (Xu et al., ; Fang et al., 2024) exhibit significantly higher time and token consumption (27.0s and 50.3k tokens, respectively) compared to EaSe (without VLM) (2.1s and 3.2k tokens). EaSe (without VLM) and ZSVG3D (Yuan et al., 2024b) have much lower costs and EaSe demonstrates a significant improvement in accuracy over ZSVG3D (Yuan et al., 2024b). Even when integrated with a VLM, EaSe maintains a more than threefold reduction in grounding time and token consumption compared to agent-based methods (Xu et al., ; Fang et al., 2024).
These quantitative results underscore the ability of EaSe to effectively balance performance and efficiency, achieving competitive accuracy while offering substantial computational cost savings compared to other training-free methods.
Table 2: Grounding time and token costs on Nr3D. EaSe has significant advantage, especially when compared to agent-based methods (VLM-Grounder and Transcrib3D). $\dagger$: evaluated on a subset having 250 samples.
Method Time/s Token
ZSVG3D 2.4 2.5k
[…]
Transcrib3D 27.0 12k
EaSe (w/o VLM) 2.1 1.2k
EaSe 7.7 (+5.6) 3.1k (+1.9k)
Table 3: Performances on ScanRefer. EaSe has close performance with VLM-Grounder and BUTD-DETR. $\dagger$: evaluated on a subset having 250 samples.
Method Acc@0.25
BUTD-DETR 52.2
[…]
Transcrib3D 51.3
EaSe 49.2
ZSVG3D 36.4
Table 4: Performances on Nr3D. VD and VID stand for view-dependent and view-independent, respectively. We only report time and token consuming for training-free methods which involve LLM or VLM. EaSe outperforms other training-free baselines in overall performance, and has significant advantages in terms of time and token consumptions. For comparison with Transcrib3D, we use GT object labels. For comparison with NS3D, we evaluate EaSe on the same subset as NS3D. $\dagger$: VLM-Grounder is evaluated on a subset having 250 samples. * : we re-run ZSVG3D using GPT-4o.
Method Overall Easy Hard VD VID
ViL3DRel 64.4 70.2 57.4 62.0 64.5
CoT3DRef 64.4 70.0 59.2 61.9 65.7
3D-VisTA 64.2 72.1 56.7 61.5 65.1
BUTD-DETR 54.6 60.7 48.4 46.0 58.0
ZSVG3D 40.2 49.1 31.1 37.8 41.6
[…]
EaSe (w/o VLM) 50.7 58.7 43.0 45.6 53.2
EaSe 52.9 - - - -
Transcrib3D 70.2 79.7 60.3 60.1 75.4
EaSe (w/o VLM) 65.7 75.6 56.2 58.7 69.1
EaSe 67.8 - - - -
NS3D 52.8 - - - -
EaSe (w/o VLM) 60.2 - - - -
4.3 Qualitative Results
Scene Visualization
Figure 3: Visualization of the grounding process. Anchor (the door) are marked with red circles. Objects that strongly match the conditions are highlighted in green, with brighter shades indicating higher matching scores.
Figure 3 illustrates a grounding process of EaSe , demonstrating how the final grounding result is constructed through the combination of multiple conditions within the referring utterance. For example, the utterance “When facing the door, it’s the shelf above the desk on the right” can be understood as following four steps, progressing from left to right in the figure. First, objects right of the door are identified using the category feature “door” and the relation feature “right”. Next, the desk right of the door is located from this set using the category feature “desk”. Subsequently, objects satisfying the joint condition above desk and on door’s right are identified. Finally, the target “shelf” is grounded from them. More visualization results can be found in the Figure 8.
4.4 Ablation Study
Figure 4: The accuracy curves of different variants. The x-axis is the generation number of the code. The y-axis is the normalized accuracy.
We conduct ablation studies to investigate the impact of various components during the code generation and optimization processes, evaluating three different variants. We choose to analyze six relations that required multiple optimization iterations. (For relatively simple relations like “small”, the generated code passes all unit tests in the first generation, so there is no optimization.)
The three variants are: Variant 1 direct prompts LLMs for multiple code, selecting the one with the highest pass rate. Variant 2 replaces the error message (see §3.3.4), which includes specific failure cases, with a general optimization instruction that omits failure cases. Variant 3 keeps all components except in-context examples. For the relations that no in-context example is used (“left”, “above”, and “corner”), variant 3 is identical to EaSe , so we only plot variant 1 and 2 on the corresponding subplots. To control for the impact of the generation results of the first iteration, we use the same results of iteration 0 across variant 2 and variant 3. In all three variants, no in-context example (§$\S$3.3.2) is used.
Figure 4 illustrates the results of the ablation study; different variants are represented by lines of different colors. The horizontal axis represents the number of iterations. The vertical axis shows the normalized accuracy on test examples associated with the relation. The effect of optimization is evident in variant 1: without optimization, LLMs fail to produce plausible spatial relation encoders for most relations, except “corner” and “between”. Variant 2 demonstrates the effect of optimization: by incorporating simple optimization, there is a noticeable performance improvement compared to variant 1. However, LLMs still struggle with relations except “above” Variant 3 highlights the effect of error messages (see § 3.3.3). By using specific failure cases in error messages, LLMs are able to generate plausible spatial relation encoders for most relations. For relations like “right”, “between” and “below” which use in-context examples, variant 3 shows a significant performance gap in initial iterations, underscoring the impact of in-context examples.
5 5. Conclusion
In this work, we propose EaSe , a training-free 3DVG method that uses Python code to encode spatial relations and an automatic pipeline for their generation and optimization. We leverage knowledge from LLMs to create spatial relation encoders, circumventing the need for human annotation or supervised learning. EaSe eliminates the need for large-scale data and offers promising advantages in accuracy and grounding cost compared to other training-free methods.
There are some limitations in EaSe . In referring utterances, there are diverse linguistic elements beyond object categories and spatial relations. Our method, focusing primarily on these core components, may struggle with non-relation words like “second from left” in complex cases. Empirically, we observe that object bounding boxes alone are insufficient for precise spatial encoding. For simplicity, we currently ignore object shapes, orientations, and different areas within the scene such as bathroom. Integrating richer scene information into LLMs and VLMs is an important avenue for future exploration.
[…]
Appendix A Prompts
In this section, we show the prompts we used.
Listing 1: Prompt for semantic parsing
You are a skilled assistant with expertise in semantic parsing.
## Task Overview
I will provide you with a sentence that describes the location of an object within a scene. Your task is to convert this description into a JSON format that captures the essential details of the object.
### The JSON object should include:
- **"category"**: string, representing the objects category.
- **"relations"**: a list of relationships between the object and other elements in the scene. Each relationship should be represented as a dictionary with the following fields:
- **"relation_name"**: string, specifying the type of relationship. The relationship can be:
- *Unary*: choose from [’corner’, on the floor’, against wall’, smaller’, larger’, taller’, lower’, within’].
- *Binary*: choose from [’above’, below’, beside’, close’, far’, left’, right’, front’, behind’, across’].
- *Ternary*: choose from [’between’, center’, middle’].
Only consider **simple** and **general** relations, donot make complex ones like "left of a blue box", "with dark appearance", "facing the window", etc. You should handle these by logical structures.
If the relationship is not mentioned in the list, you should choose the most appropriate relation above. **Never** create a new relation name!
- **"objects"**: a list of objects involved in the relationship. Every object in the list should have the same JSON structure. The list structure depends on the relationship type:
- *Unary*: The list should be empty.
- *Binary*: The list should contain one object.
- *Ternary*: The list should contain two objects.
- **"negative"**: boolean, indicating if the object is explicitly described as not having this relationship. Set this to True if applicable.
## Guidelines:
- First, generate a plan outlining the objects appearance and relationships based on the sentence. Then, use this plan to create the JSON representation.
## Examples:
### Example 1:
**Sentence**: The correct whiteboard is the one on a table.
**Plan**: "Correct" does not describe appearance. The appearances are "whiteboard" and "table", and the "whiteboard" is on the "table".
**Parsed JSON**:
“‘json
"category": "whiteboard",
"relations": [
"relation_name": "above",
"objects": [
"category": "table",
"relations": []
2 more examples.
Listing 2: An example of initial prompt for relation encoder generation
You are an expert on spatial relation analysis and code generation.
# Introduction to task
Your task is to write a Python class which can be used to compute the metric value of the existence of some spatial relationship between two objects in a 3D scene given their positions and sizes. Higher the metric value, higher the probability of the two objects have that relation.
In the class, you will be given the positions and sizes of the objects in the scene. The class should have a method forward which returns a tensor of shape (N, N), where element (i, j) is the metric value of the relation between object i and object j.
In the 3D scene, x-y plane is the horizontal plane, z-axis is the vertical axis.
# Introduction to programming environment
Here is an example class for Left relation. The class you write should have the same structure as the example class.
“‘python
class Left:
Make sure all tensors are placed on DEVICE‘, which has been defined in the environment.
The code output should be formatted as a python code string: "“‘python “‘".
# Some helpful tips
(1) You should only use the given variables, and you should not introduce new variables.
(2) The metric value should be sensitive to the input arguments, which means if the arguments change a little, the value should change a lot.
(3) The metric value should be 0 if the two objects dont have that relation, never set negative values!
(4) Never treat an object as its center point, you must consider the size of the bounding box, just like the example code. Never set an threshold to determine the relation. The value of the relation should be continuous and sparse.
(5) You should imagine that you are at position (0, 0) to determine the relative positions.
(6) Remember you are **in** the scene and look around, not look from the top. So never use the same way as 2D environment.
Propose your method first and then generate the code. Think step by step.
Dont use any axis or specific direction as the reference direction or right direction, your method should work for any perspectives.
Listing 3: An example for feedback message
We have run your code on some cases. Here are 3 failure cases:
# Case 1.
Metric value of object tensor([ 0.3992, -0.5619, 0.8831, 0.3921, 0.3476, 0.1059], device=’mps:0’) "above" object tensor([-0.0432, -0.6965, 0.8483, 0.6526, 0.4943, 0.3061], device=’mps:0’) should be larger than 0. Metric value of object tensor([ 0.3992, -0.5619, 0.8831, 0.3921, 0.3476, 0.1059], device=’mps:0’) "above" object tensor([-0.0432, -0.6965, 0.8483, 0.6526, 0.4943, 0.3061], device=’mps:0’) should be higher than the metric value of object tensor([0.5338, 1.1607, 1.1160, 0.2121, 0.3323, 0.8192], device=’mps:0’) "above" object tensor([-0.0432, -0.6965, 0.8483, 0.6526, 0.4943, 0.3061], device=’mps:0’).
more 2 cases
The first three are the center of the object, the last three are the size of the object. x-y is the horizontal plane and z is the vertical axis.
After test, the pass rate of your code is too low. So you MUST check carefully where the problem is. If you cant find the problem, you should
come up with a new algorithm and re-write your code.
Dont forget the following tips:
(1) You should imagine that you are at position (0, 0, 0) to determine the relative positions.
(2) Remember you are **in** the scene and look around, not look from the top. So never use the same way as 2D environment.
(3) Dont use any of x-axis or y-axis as your perspective, Your method should work for every perspective.
(4) The horizontal plane is x-y plane.
Please carefully analyze each of the failure case and explain why your code failed to pass it. The reason can be incorrect test case might or your code might not be able to handle some specific cases. Please write your analysis for each of the failure cases.
After the analysis of all cases, you should write the improved code based on your analysis. But **never** modify on the class methods and function parameters.
Some possible improvement ways:
1. Use a new algorithm to calculate the metric value rather than just modifying the existing code.
2. Consider carefully what other factors might be relevant to the spatial relationship between two objects and use them in your calculation.
3. Check the correctness of the input data and the calculation process.
Listing 4: The general prompt for self-reflection in variant 2 of ablation studies.
Reflect on the code above, think carefully how to make it better. For example, check if you ignore some factors that may affect the result or use a wrong method.
Then you must re-write the code in the same format. Remeber all the tips!
Appendix B Implementation Details
B.1 Definition of Symbolic Expression
The semantic parser translates the referring utterance $\mathcal{U}$ into a JSON-based symbolic expression $\mathcal{E}$ comprising the following elements:
  • Category: A string indicating the category of the target object referenced in $\mathcal{U}$.
  • Relations: A list specifying spatial constraints relative to the target object. Each entry in this list contains:
    • relation_name: A string identifying the spatial relation in $\mathcal{U}$ (e.g., “near,” “above”).
    • anchors: A list of objects that share the given spatial relation with the target object. Each object is represented as its own JSON entity.
    • negative: A boolean value which, if set to true, denotes that the target object should not exhibit the specified spatial relation.
Appendix C Definition of Symbolic Expression
C.1 Example of Relation Encoder
An example of relation encoder for relation “Near" is shown in Figure 5. The class is initialized using the bounding boxes of all objects in the scene. he function _init_params computes the distances between all object pairs. The forward function calculates the relation feature for "Near" by performing numerical operations on these distances.
class Near:
def __init__(
self,
object_locations: torch.Tensor) -> None:
Args:
object_locations: torch.Tensor, shape (N, 6),
N is the number of objects in the scene.
self.object_locations = object_locations.to(DEVICE)
self._init_params()
def _init_params(self) -> None:
Computing some necessary parameters about ‘Near‘ relation and
initialize ‘self.param‘.
# Based on average size to calculate a baseline distance
sizes = self.object_locations[:, 3:]
self.avg_size_norm = sizes.mean(dim=0).norm().to(DEVICE)
def forward(self) -> torch.Tensor:
Return a tensor of shape (N, N),
where element (i, j) is the metric value
of the ‘Near‘ relation between object
i and object j.
centers = self.object_locations[:, :3]
sizes = self.object_locations[:, 3:]
# Calculate the pairwise distances between object
centers in a vectorized manner
diff = centers.unsqueeze(1) - centers.unsqueeze(0)
distances = diff.norm(dim=2)
# Calculate the "nearness" metric based
on distances and average size norm
nearness_metric = torch.exp(-distances /
(self.avg_size_norm + 1e-6))
# Set diagonal to zero since an object
cannot be near itself
nearness_metric.fill_diagonal_(0)
return nearness_metric.to(DEVICE)
Figure 5: Example relation encoder for “Near".
C.1.1 In Context Example Selection
The graph used for select in the context example for relation encoder generation is shown in Figure 6.
Figure 6: The graph for in context example selection. An edge from node A to node B means that the encoder for relation A is used as an in context example when generating for relation B.
Category Features
Yuan et al. (2024b) provides the predicted classified label for each object. For the category feature $f_{\text{category}}\in\mathbb{R}^{N}$ ($N$ is the number of objects), we compute the cosine similarity $sim\in\mathbb{R}^{N}$ between the category and the predicted labels using CLIP (Radford et al., 2021). Subsequently, we define the category feature as:
$f_{\text{category}}=softmax(100\times sim)$
To optimize computation, we cache features for different categories in a dictionary $C$.
Relation Features
We use the object bounding boxes in the scene to initialize the relation encoders and then call the forward() function to compute the corresponding relation feature, f_rel. These relation features are also cached in a dictionary $R$ for efficient reuse.
Execution Details
[…]
Appendix D More Quantitative Results
NS3D
We show evaluation results in NS3D(Hsu et al., 2023) in Table 4 . NS3D can only learn concepts (e.g. relation name, category name) from the training set and its parsing results of Nr3D contain more than 5,000 concepts, resulting in a long-tailed problem. So it selects a subset containing 1,041 test examples, which only contains the same concepts as Sr3D, the dataset it is mainly trained on. On the NS3D subset, EaSe has a large improvement over Ns3D (more than 7%), which shows the advantage of EaSe for processing natural grounding tasks.
Sr3D
In Table 5, we show the evaluation results on NS3D, a subset of Nr3D, proposed by Hsu et al. (2023). Even not using training data of Sr3D, EaSe still achieves comparable performance with Hsu et al. (2023) on both settings (w/ and w/o GT labels). We show evaluation results on Sr3D, a subset of Achlioptas et al. (2020) in Table 5. In the standard setting, if using GT object labels, the accuracy of our method (w/o VLM) on Sr3D is 95.3%, the performance of NS3D and Fang et al. (2024) are 96.9% and 98.4%. So we believe that the bottleneck of Sr3D performance is object detection and classification rather than spatial relation understanding because its relation annotation is synthesized by relatively simple functions. So we mainly focus on natural benchmarks (Nr3D and Scanrefer) which have complex and real spatial relations.
Table 5: Performance on Sr3D.
Method Overall NS3D
BUTD-DETR 67.0 -
NS3D (Hsu et al., 2023) 62.7 52.6
NS3D (Hsu et al., 2023) (w/ GT Label) 96.9 -
Transcrib3D (Fang et al., 2024) (w/ GT Label) 98.4 -
Ours (w/o VLM) 62.0 58.6
Ours (w/o VLM, w/ GT Label) 95.1 -
Appendix E More Qualitative Results
E.1 Effect of Code Optimization
We show the change between the initial response and the final code after multiple rounds of sampling and iterative refinement. By transitioning from a strict geometric overlap calculation to a continuous, exponential-based measure for both vertical and horizontal distances, the optimized code now captures nuances in “above” relation more robustly. This improved formulation inherently handles scenarios where objects are close but not strictly overlapping, and it provides a more stable gradient for training. Consequently, the updated model passes all test cases by offering a smoother, more differentiable metric that better aligns with real-world spatial relations and passes more test cases.
class Above:
def __init__(self, object_locations: torch.Tensor) -> None:
Args:
object_locations: torch.Tensor, shape (N, 6),
N is the number of objects in the scene.
The first three columns are the center of the object (x, y, z),
the last three columns are the size of the object (width, height, depth).
self.object_locations = object_locations.to(DEVICE)
self._init_params()
def _init_params(self) -> None:
Compute necessary parameters about ‘OnTopOf‘
relation and initialize self.param if needed.
# shape (N, 3)
self.centers = self.object_locations[:, :3]
# shape (N, 3)
self.sizes = self.object_locations[:, 3:]
def forward(self) -> torch.Tensor:
Return a tensor of shape (N, N), where element (i, j) is the
metric value of the ‘OnTopOf‘ relation between object i and object j.
N = self.object_locations.shape[0]
# Expand dimensions for broadcasting
# Shape (N, 1, 3)
centers_i = self.centers.unsqueeze(1)
# Shape (N, 1, 3)
sizes_i = self.sizes.unsqueeze(1)
# Shape (1, N, 3)
centers_j = self.centers.unsqueeze(0)
# Shape (1, N, 3)
sizes_j = self.sizes.unsqueeze(0)
# Calculate vertical distances
vertical_distance = centers_i[..., 2] - centers_j[..., 2]
- (sizes_i[..., 1] + sizes_j[..., 1]) / 2
# Calculate horizontal overlaps
overlap_x = torch.clamp((sizes_i[..., 0] + sizes_j[..., 0]) / 2
- torch.abs(centers_i[..., 0] - centers_j[..., 0]), min=0)
overlap_y = torch.clamp((sizes_i[..., 2] + sizes_j[..., 2]) / 2
- torch.abs(centers_i[..., 1] - centers_j[..., 1]), min=0)
# Calculate the overlap area in the x-y plane
horizontal_overlap_area = overlap_x * overlap_y
# Calculate the metric value, only where object i is above object j
metric_value = torch.where(vertical_distance > 0,
horizontal_overlap_area / (1 + vertical_distance),
torch.tensor(0.0, device=DEVICE))
# Mask out the diagonal (i == j) to ensure
self-relations are not considered
metric_value.fill_diagonal_(0)
return metric_value
(a) Initial generation of relation encoder of “above".
class Above:
def __init__(
self,
object_locations: torch.Tensor) -> None:
Args:
object_locations: torch.Tensor, shape (N, 6),
N is the number of objects in the scene.
The first three columns are the center of the object (x, y, z),
the last three columns are the size of the object (width, height, depth).
self.object_locations = object_locations
self._init_params()
def _init_params(self) -> None:
Compute necessary parameters about ‘OnTopOf‘
relation and initialize parameters.
self.center = self.object_locations[:, :3] # (x, y, z)
self.size = self.object_locations[:, 3:] # (width, depth, height)
def forward(self) -> torch.Tensor:
Return a tensor of shape (N, N), where element (i, j)
is the metric value of the ‘OnTopOf‘
relation between object i and object j.
N = self.object_locations.shape[0]
# Calculate bottom of i and top of j
height_i = self.size[:, 2].view(N, 1)
bottom_of_i = self.center[:, 2].view(N, 1) - height_i / 2
top_of_j = self.center[:, 2].view(1, N)
+ self.size[:, 2].view(1, N) / 2
# Vertical proximity
vertical_proximity = torch.exp(-torch.abs(bottom_of_i - top_of_j)
/ (height_i / 2))
# Center distances
center_dist_x = torch.abs(self.center[:, 0].view(N, 1)
- self.center[:, 0].view(1, N))
center_dist_y = torch.abs(self.center[:, 1].view(N, 1)
- self.center[:, 1].view(1, N))
# Combined sizes
combined_size_x = (self.size[:, 0].view(N, 1) + self.size[:, 0].view(1, N)) / 2
combined_size_y = (self.size[:, 1].view(N, 1) + self.size[:, 1].view(1, N)) / 2
# Horizontal alignment
horizontal_alignment = torch.exp(-(center_dist_x / combined_size_x) - (center_dist_y / combined_size_y))
# Combine the metrics with a weight emphasizing more on vertical proximity
relation_metric = vertical_proximity * horizontal_alignment
# Remove self-relations by zeroing the diagonal
relation_metric.fill_diagonal_(0)
return relation_metric
(b) The relation encoder of “above" after multi-turn sampling and optimization
Figure 7: Example of code optimization result on relation encoder of “above". The left output is from the first iteration, which only pass 18 test cases out of 37. After 3 iterations of sampling and optimization, we get the the relation encoder as the right one. It passes all 37 test cases.
E.2 Relation Constraints
Feng et al. (2024) propose that certain spatial relations are symmetric, like “near" or “far", meaning that if object A is "near" B, then B should also be "near" A. Consequently, the features of these relations should be symmetric. Conversely, other relations are inherently asymmetric, such as “left" or “right". For these relations, if a feature element is positive (indicating the presence of the spatial relation), its corresponding symmetric element should be negative (indicating the absence of the reverse spatial relation).
Feng et al. (2024) leverage large language models (LLMs) to annotate these constraints and apply an auxiliary loss to enforce them during training. In contrast, while EaSe does not explicitly train or use specialized instructions to create these constraints, we observe that some LLM-generated relation encoders inherently produce relation features that satisfy these constraints. Moreover, for certain relations, these constraints are guaranteed due to the deterministic execution of the code. In Figure 10, we present four relation features for a scene:
  • Features for “near" and “far" are guaranteed to be symmetric.
  • For asymmetric features such as “left" and “right," if $f_{i,j}>0$, it is guaranteed that $f_{j,i}=0$.
E.3 Condition Level Accuracy
Our parsed symbolic expressions typically include one or more spatial conditions for the target object. However, some conditions in the referring utterance may be redundant.
For instance, if the referring utterance is "find the monitor on the floor and under the desk," and all monitors "on the floor" are also "under the desk," then one of these two conditions is redundant. This means that even if the method fails to process one of the conditions, it can still provide the correct grounding result.
To better understand EaSe ’s capability, we evaluate it on utterances containing a single condition.
We categorize objects of the same class into groups. Within each group, we collect the conditions for each object from the parsed expressions. Each condition is represented in JSON format, such as: "relation": ..., "anchors": [...].
These conditions are executed seamlessly to identify the best-matching object. We compute the average precision and recall for all condition-level matches. EaSe achieves an average precision of 67.5% and an average recall of 66.9%.
Require : symbolic expression $E$, category features $C$, relation features $R$
1
Output : matching_score
1
[…]
6
[…]
8
[…]
10
[…]
12
[…]
14
15
[…]
17             $\texttt{a}\leftarrow\texttt{Execute}(\texttt{anchors}[0])$
18
[…]
20
[…]
23
[…]
25
[…]
27
[…]
29
30
[…]
32
[…]
34
35            $\texttt{f}\leftarrow\texttt{max}(\texttt{f})-\texttt{f}$
36
[…]
38
39
Output : matching_score
Algorithm 2 Execute
E.4 More Visualization Results
We visualize more grounding examples in LABEL:fig:appendix. The first row illustrates the grounding process for the kitchen cabinet close to the fridge and beside the stove. In the process, the stove, objects beside the stove, and the objects near the fridge are sequentially grounded, culminating in the target kitchen cabinet highlighted in green. The second row shows the grounding process for right trash can below the sink. Starting with the objects below the sink, followed by the objects on the right of the sink, and finally combining these conditions to highlight the target trash can in green.
E.5 Effect of Unit tests
To demonstrate the impact of filtering generated code based on its accuracy on training cases, we selected six relations and plotted their performance. The x-axis represents the pass rate on training cases, while the y-axis shows the number of passed examples in all relevant test cases.
For straightforward relations such as “near" or “far", GPT-4o can pass all unit tests on the first attempt, so we focus on cases requiring multiple refinement steps.
The results, shown in Figure 9, indicate that for five out of six relations (excluding behind), the code with the highest pass rate on training cases achieves top-tier performance on the test set. However, for the behind relation, using the best-performing code on the training cases results in about 15 fewer passed test cases compared to using code with approximately 70% accuracy. Despite this, it still outperforms code with accuracy below 0.5.
This discrepancy for behind may stem from biases in the training data collection process. Overall, selecting code based on its performance on the training set is effective for achieving strong test set performance.
Figure 8: The target objects are: “Stove next to another stove and close to the fridge" (top row) and “Trashcan to the right of and below the sink" (bottom row).
Figure 9: Corresponding relation between the unit test pass rate and number of correct examples on test set.
Figure 10: Relation feature examples. The features of “near" and “far" are symmetric, meaning mutual relationships hold true in both directions. For “left" and “right," if an element is positive, its corresponding symmetric element is zero, ensuring asymmetry. Additionally, all diagonal elements are zero, as self-relations are not considered.